OpenStudy (anonymous):

Simplify. Not sure how to simplify this anybody??? Looked on YouTube but nothing similar please explain

OpenStudy (blackfireskull):

what is the question you are trying to simplify?

OpenStudy (anonymous):

Not sure how to type it I'll take a picture

OpenStudy (saurabh135):

see (2)^2 means 2*2 similarly (5/6)^2= (5/6)*(5/6)=25/36

OpenStudy (anonymous):

Thank youuu I get it and I finally know how to write the problem😊